# On the Lipschitz Continuity of Set Aggregation Functions and Neural Networks for Sets

### v2

@@ Appendix C Experimental Setup
C.1 Lipschitz Constant of Aggregation Functions
@@
We represent each document as a multiset of the embeddings of its words. The word embeddings are obtained from a pre-trained model which contains $300$-dimensional vectors [14]. The embeddings are first fed to an MLP, and then an aggregation function (i. e., sum, mean or max) is applied which produces a single vector representation for each document. This vector representation is passed onto another MLP. The two MLPs consist of two layers and the ReLU function is applied to the output of the first layer of each MLP. The output of the second MLP is followed by the softmax function. The hidden dimension size of all layers is set to $64$. The model is trained by minimizing the cross-entropy loss. The minimization is performed using Adam with a learning rate equal to $0.001$. The number of epochs is set to $200$ and the batch size to $64$. At the end of each epoch, we compute the performance of the model on the validation set, and we choose as our final model the one that achieved the smallest loss on the validation set. As discussed above, in each experiment, we compute the distance of the “latent” multisets that emerge at the output of the first MLP (just before the aggregation function is applied) of the model that achieves the lowest validation loss.
C.2 Lipschitz Constant of Neural Networks for Sets
@@
We represent each document as a multiset of the embeddings of its words. The word embeddings are obtained from a publicly available pre-trained model [14]. We randomly split the dataset into training, validation, and test sets with a $80:10:10$ split ratio. The embeddings are first fed to a fully-connected layer, and then an aggregation function (i. e., sum, mean or max) is applied which produces a single vector representation for each document. This vector representation is passed onto another fully-connected layer. The ReLU function is applied to the emerging vector and then a final fully-connected layer followed by the softmax function outputs class probabilities. The output dimension of the first two fully-connected layers is set to $64$. The model is trained by minimizing the cross-entropy loss. The minimization is performed using Adam with a learning rate equal to $0.001$. The number of epochs is set to $200$ and the batch size to $64$. At the end of each epoch, we compute the performance of the model on the validation set, and we choose as our final model the one that achieved the smallest loss on the validation set. As discussed above, we can exactly compute the Lipschitz constants of the two fully-connected layers.
C.3 Stability under Perturbations of Input Multisets
@@
Each document of the Polarity dataset is represented as a multiset of word vectors. The word vectors are obtained from a publicly available pre-trained model [14]. We randomly split the dataset into training, validation, and test sets with a $80:10:10$ split ratio. The $\textsc{NN}_{\textsc{mean}}$ and $\textsc{NN}_{\textsc{max}}$ models consist of an MLP which transforms the representations of the words, an aggregation function (mean and max, respectively) and a second MLP which produces the output. Both MLPs consist of two hidden layers. The ReLU function is applied to the outputs of the first layer. The hidden dimension size is set to $64$ for all hidden layers. The model is trained for $20$ epochs by minimizing the cross-entropy loss function with the Adam optimizer and a learning rate of $0.001$. At the end of each epoch, we compute the performance of the model on the validation set, and we choose as our final model the one that achieved the smallest loss on the validation set.
C.4 Generalization under Distribution Shifts
This set of experiments is conducted on the Polarity dataset. The architecture of $\textsc{NN}_{\textsc{mean}}$ and $\textsc{NN}_{\textsc{max}}$, and the training details are same as in subsection C.3 above.
Figure 4: Each dot corresponds to a textual document from the test set of Polarity that is represented as a multiset of word vectors. Each subfigure compares the distance of documents computed by a distance function for multisets (i. e., EMD, Hausdorff distance or matching distance) against the Euclidean distance of the representations of the documents that emerge after the application of an aggregation function (i. e., mean, sum or max). The correlation between the two distances is also computed and visualized. The Lipschitz bounds (dash lines) upper bound the distances of the outputs of the aggregation functions.
Appendix D Additional Results
D.1 Lipschitz Constant of Aggregation Functions
@@
Figure 5: Each dot corresponds to a textual document from the test set of Polarity. Each subfigure compares the distance of input documents represented as multisets of word vectors computed by EMD, Hausdorff distance or matching distance against the Euclidean distance of the representations of the documents that emerge at the second-to-last layer of $\textsc{NN}_{\textsc{mean}}$, $\textsc{NN}_{\textsc{sum}}$ or $\textsc{NN}_{\textsc{max}}$. The correlation between the two distances is also computed and reported. The Lipschitz bounds (dash lines) upper bound the distances of the outputs of the neural networks.
D.2 Upper Bounds of Lipschitz Constants of Neural Networks for Sets
@@
We next provide some additional empirical results that validate the findings of Theorem 3.1 (i. e., upper bounds of Lipschitz constants of neural networks for sets). Since the Polarity dataset consists of documents which might differ from each other in the number of terms, by Theorem 3.3 we can derive upper bounds only for $2$ out of the $9$ combinations of distance functions for multisets and aggregation functions. While the dash lines upper bound the Euclidean distance of the outputs of the aggregation functions, both bounds are relatively loose on the Polarity dataset. The correlations between the distances of the representations produced by the neural network for the multisets and the distances produced by distance functions for multisets are much lower than those of the previous experiments. The highest correlation is equal to $0.55$ (between the neural network that utilizes the mean aggregation function and EMD), while there are even negative correlations. This is not surprising since for most of the combinations, there are no upper bounds on the Lipschitz constant of the corresponding neural networks.
@@
D.3 Stability of Neural Networks for Sets under Perturbations
Here we provide some further details about the experiments presented in subsection 3.3. Specifically, for the experiments conducted on ModelNet40, we attribute the drop in performance of $\textsc{NN}_{\textsc{max}}$ to the large Hausdorff distances between each test sample and its perturbed version. For each test sample $X_{i}$ (where $i\in[2468]$), let $X_{i}^{\prime}$ denote the multiset that emerges from the application of Pert. #1 to $X_{i}$. Let also $y_{i}$ denote the class label of sample $X_{i}$. We compute the Hausdorff distance between $X_{i}$ and $X_{i}^{\prime}$ ($d_{H}(X_{i},X_{i}^{\prime})$). We then compute the average Hausdorff distance between $X_{i}$ and the rest of the test samples that belong to the same class as $X_{i}$. Let $\mathcal{S}_{i}=\{X_{j}\colon j\in[2468],y_{i}=y_{j}\}$ denote the set of all test samples that belong to the same class as $X_{i}$. Then, we compute $\bar{d}_{H}(X_{i},\mathcal{S}_{i})$ as follows:
We then compute $d_{H}(X_{i},X_{i}^{\prime})-\bar{d}_{H}(X_{i},\mathcal{S}_{i})$. If this value is positive, then the distance from $X_{i}$ to $X_{i}^{\prime}$ is greater than the average distance of $X_{i}$ to the other multisets that belong to the same class as $X_{i}$. We calculated this value for all test samples and then computed the average value which was found to be $2.63(\pm 1.10)$. In general, the perturbation increases the upper bound of the Lipschitz constant of $\textsc{NN}_{\textsc{max}}$ compared to the upper bound for samples that belong to the same class, and thus $X_{i}$ and $X_{i}^{\prime}$ might end up having dissimilar representations. On the other hand, for EMD, the average distance is equal to $-1.18(\pm 0.72)$. For EMD the upper bound is in general tighter than the bound for pairs of multisets that belong to the same class, and this explains why $\textsc{NN}_{\textsc{mean}}$ is robust to Pert. #1.
D.4 Generalization under Distribution Shifts
To evaluate the generalization of the two Lipschitz continuous models ($\textsc{NN}_{\textsc{mean}}$ and $\textsc{NN}_{\textsc{max}}$) under distribution shifts, we also experiment with the Amazon review dataset [5]. The dataset consists of product reviews from Amazon for four different types of products (domains), namely books, DVDs, electronics and kitchen appliances. For each domain, there exist $2,000$ labeled reviews (positive or negative) and the classes are balanced. We construct $4$ adaptation tasks. In each task, the $\textsc{NN}_{\textsc{mean}}$ and $\textsc{NN}_{\textsc{max}}$ models are trained on reviews for a single type of products and evaluated on all domains.
@@ D.4 Generalization under Distribution Shifts
Each review is represented as a multiset of word vectors. The word vectors are obtained from a publicly available pre-trained model [14]. The $\textsc{NN}_{\textsc{mean}}$ and $\textsc{NN}_{\textsc{max}}$ models consist of an MLP which transforms the representations of the words, an aggregation function (mean and max, respectively) and a second MLP which produces the output. Both MLPs consist of two hidden layers. The ReLU function is applied to the outputs of the first layer and also dropout is applied between the two layers with $p=0.2$. The hidden dimension size is set to $64$ for all hidden layers. The model is trained for $50$ epochs by minimizing the cross-entropy loss function with the Adam optimizer and a learning rate of $0.001$. At the end of each epoch, we compute the performance of the model on the validation set, and we choose as our final model the one that achieved the smallest loss on the validation set.
@@
Figure 6 illustrates the Wasserstein distance with $p=1$ between groups when the EMD (Top Right) and the Hausdorff distance (Bottom Right) are used as ground metrics. It also shows the drop in accuracy when the $\textsc{NN}_{\textsc{mean}}$ (Top Left) and $\textsc{NN}_{\textsc{max}}$ (Bottom Left) models are trained on one domain and evaluated on the others. Each row corresponds to one specific model, e. g., the first row represents the model trained on the reviews for books. We observe that in general the drop in accuracy follows a similar pattern with the distance between the source and target domains, i. e., the higher the distance the higher the drop in accuracy. We also computed the Pearson correlation between the drop in accuracy and the domain dicrepancies. We found that the Wasserstein distance based on EMD highly correlates with the accuracy drop of $\textsc{NN}_{\textsc{mean}}$ ($r=0.917$), while there is an even higher correlation between the Wasserstein distance based on Hausdorff distance and the accuracy drop of $\textsc{NN}_{\textsc{max}}$ ($r=0.941$).

### v3

@@ Appendix D Experimental Setup
D.1 Lipschitz Constant of Aggregation Functions
@@
We represent each document as a multiset of the embeddings of its words. The word embeddings are obtained from a pre-trained model which contains $300$-dimensional vectors (Mikolov et al., 2013). The embeddings are first fed to an MLP, and then an aggregation function (i. e., sum, mean or max) is applied which produces a single vector representation for each document. This vector representation is passed onto another MLP. The two MLPs consist of two layers and the ReLU function is applied to the output of the first layer of each MLP. The output of the second MLP is followed by the softmax function. The hidden dimension size of all layers is set to $64$. The model is trained by minimizing the cross-entropy loss. The minimization is performed using Adam with a learning rate equal to $0.001$. The number of epochs is set to $200$ and the batch size to $64$. At the end of each epoch, we compute the performance of the model on the validation set, and we choose as our final model the one that achieved the smallest loss on the validation set. As discussed above, in each experiment, we compute the distance of the “latent” multisets that emerge at the output of the first MLP (just before the aggregation function is applied) of the model that achieves the lowest validation loss.
D.2 Lipschitz Constant of Neural Networks for Sets
@@
We represent each document as a multiset of the embeddings of its words. The word embeddings are obtained from a publicly available pre-trained model (Mikolov et al., 2013). We randomly split the dataset into training, validation, and test sets with a $80:10:10$ split ratio. The embeddings are first fed to a fully-connected layer, and then an aggregation function (i. e., sum, mean or max) is applied which produces a single vector representation for each document. This vector representation is passed onto another fully-connected layer. The ReLU function is applied to the emerging vector and then a final fully-connected layer followed by the softmax function outputs class probabilities. The output dimension of the first two fully-connected layers is set to $64$. The model is trained by minimizing the cross-entropy loss. The minimization is performed using Adam with a learning rate equal to $0.001$. The number of epochs is set to $200$ and the batch size to $64$. At the end of each epoch, we compute the performance of the model on the validation set, and we choose as our final model the one that achieved the smallest loss on the validation set. As discussed above, we can exactly compute the Lipschitz constants of the two fully-connected layers.
D.3 Stability under Perturbations of Input Multisets
@@
Each document of the Polarity dataset is represented as a multiset of word vectors. The word vectors are obtained from a publicly available pre-trained model (Mikolov et al., 2013). We randomly split the dataset into training, validation, and test sets with a $80:10:10$ split ratio. The $\textsc{NN}_{\textsc{mean}}$ and $\textsc{NN}_{\textsc{max}}$ models consist of an MLP which transforms the representations of the words, an aggregation function (mean and max, respectively) and a second MLP which produces the output. Both MLPs consist of two hidden layers. The ReLU function is applied to the outputs of the first layer. The hidden dimension size is set to $64$ for all hidden layers. The model is trained for $20$ epochs by minimizing the cross-entropy loss function with the Adam optimizer and a learning rate of $0.001$. At the end of each epoch, we compute the performance of the model on the validation set, and we choose as our final model the one that achieved the smallest loss on the validation set.
D.4 Generalization under Distribution Shifts
This set of experiments is conducted on the Polarity dataset. The architecture of $\textsc{NN}_{\textsc{mean}}$ and $\textsc{NN}_{\textsc{max}}$, and the training details are same as in subsection D.3 above.
Figure 4: Each dot corresponds to a pair of documents from the test set of Polarity that is represented as a multiset of word vectors. Each subfigure compares the distance between the latent representations of pairs of documents computed by a distance function for multisets (i. e., EMD, Hausdorff distance or matching distance) with the Euclidean distance between the representations of the pairs obtained after applying an aggregation function (i. e., mean, sum or max). The correlation between the two distances is also computed and visualized. The Lipschitz bounds are illustrated with dashed lines.
Appendix E Additional Results
E.1 Lipschitz Constant of Aggregation Functions
@@
Figure 5: Each dot corresponds to a pair of documents from the test set of Polarity that is represented as a multiset of word vectors. Each subfigure compares the distance between the pairs of documents computed by a distance function for multisets (i. e., EMD, Hausdorff distance or matching distance) with the Euclidean distance between the representations of the pairs that emerge at the second-to-last layer of $\textsc{NN}_{\textsc{mean}}$, $\textsc{NN}_{\textsc{sum}}$ or $\textsc{NN}_{\textsc{max}}$. The correlation between the two distances is also computed and visualized. The Lipschitz bounds are illustrated with dashed lines.
E.2 Upper Bounds of Lipschitz Constants of Neural Networks for Sets
@@
We next provide some additional empirical results that validate the findings of Theorem 3.4 (i. e., upper bounds of Lipschitz constants of neural networks for sets). The results are shown in Figure 5. Since the Polarity dataset consists of documents which might differ from each other in the number of terms, by Theorem 3.4 we can derive upper bounds only for $2$ out of the $9$ combinations of distance functions for multisets and aggregation functions. While the dash lines upper bound the Euclidean distance of the outputs of the aggregation functions, both bounds are relatively loose on the Polarity dataset. The correlations between the distances of the representations produced by the neural network for the multisets and the distances produced by distance functions for multisets are much lower than those of the previous experiments. The highest correlation is equal to $0.55$ (between the neural network that utilizes the mean aggregation function and EMD), while there are even negative correlations. This is not surprising since for most of the combinations, there are no upper bounds on the Lipschitz constant of the corresponding neural networks.
Figure 6: Each dot corresponds to a pair of point clouds from the test set of ModelNet40. Each subfigure compares the distance between the pairs of point clouds computed by EMD, Hausdorff distance or matching distance with the Euclidean distance between the representations of the pairs that emerge at the second-to-last layer of a neural network that consists of a fully-connected layer, an attention mechanism that aggregates the representations of the elements and two more fully-connected layers. The correlation between the two distances is also computed and visualized.
We also visualize the relationship between the output of the three considered distance functions and the Euclidean distance of the multiset representations that are produced by a neural network that utilizes the attention mechanism that is presented in subsection 3.1. Specifically, the neural network is identical to $\textsc{NN}_{\textsc{mean}}$, $\textsc{NN}_{\textsc{sum}}$ and $\textsc{NN}_{\textsc{max}}$, but instead of the standard aggregation functions, it employs the aforementioned attention mechanism. The experiments are conducted on the ModelNet40 dataset and the results are shown in Figure 6. We have shown that the attention mechanism is not Lipschitz continuous with respect to any of the three considered distance functions, and therefore the neural network models that employ this mechanism are also not Lipschitz continuous. We observe in Figure 6 that the correlations are indeed much lower than those illustrated in Figure 2 which confirms our theoretical result.
E.3 Stability of Neural Networks for Sets under Perturbations
Here we provide some further details about the experiments presented in subsection 4.3. Specifically, for the experiments conducted on ModelNet40, we attribute the drop in performance of $\textsc{NN}_{\textsc{max}}$ to the large Hausdorff distances between each test sample and its perturbed version. For each test sample $X_{i}$ (where $i\in[2468]$), let $X_{i}^{\prime}$ denote the multiset that emerges from the application of Pert. #1 to $X_{i}$. Let also $y_{i}$ denote the class label of sample $X_{i}$. We compute the Hausdorff distance between $X_{i}$ and $X_{i}^{\prime}$ (i. e., $d_{H}(X_{i},X_{i}^{\prime})$). We then compute the average Hausdorff distance between $X_{i}$ and the rest of the test samples that belong to the same class as $X_{i}$. Let $\mathcal{S}_{i}=\{X_{j}\colon j\in[2468],y_{i}=y_{j}\}$ denote the set of all test samples that belong to the same class as $X_{i}$. Then, we compute $\bar{d}_{H}(X_{i},\mathcal{S}_{i})$ as follows:
We then compute $d_{H}(X_{i},X_{i}^{\prime})-\bar{d}_{H}(X_{i},\mathcal{S}_{i})$. If this value is positive, then the distance from $X_{i}$ to $X_{i}^{\prime}$ is greater than the average distance of $X_{i}$ to the other multisets that belong to the same class as $X_{i}$. We calculated this value for all test samples and then computed the average value which was found to be $2.63(\pm 1.10)$. In general, the perturbation increases the upper bound of the Lipschitz constant of $\textsc{NN}_{\textsc{max}}$ compared to the upper bound for samples that belong to the same class, and thus $X_{i}$ and $X_{i}^{\prime}$ might end up having dissimilar representations. On the other hand, for EMD, the average distance is equal to $1.18(\pm 0.72)$. For EMD the upper bound is in general tighter than the bound for pairs of multisets that belong to the same class, and this explains why $\textsc{NN}_{\textsc{mean}}$ is robust to Pert. #1.
E.4 Generalization under Distribution Shifts
To evaluate the generalization of the two Lipschitz continuous models ($\textsc{NN}_{\textsc{mean}}$ and $\textsc{NN}_{\textsc{max}}$) under distribution shifts, we also experiment with the Amazon review dataset (Blitzer et al., 2007). The dataset consists of product reviews from Amazon for four different types of products (domains), namely books, DVDs, electronics and kitchen appliances. For each domain, there exist $2,000$ labeled reviews (positive or negative) and the classes are balanced. We construct $4$ adaptation tasks. In each task, the $\textsc{NN}_{\textsc{mean}}$ and $\textsc{NN}_{\textsc{max}}$ models are trained on reviews for a single type of products and evaluated on all domains.
@@ E.4 Generalization under Distribution Shifts
Each review is represented as a multiset of word vectors. The word vectors are obtained from a publicly available pre-trained model (Mikolov et al., 2013). The $\textsc{NN}_{\textsc{mean}}$ and $\textsc{NN}_{\textsc{max}}$ models consist of an MLP which transforms the representations of the words, an aggregation function (mean and max, respectively) and a second MLP which produces the output. Both MLPs consist of two hidden layers. The ReLU function is applied to the outputs of the first layer and also dropout is applied between the two layers with $p=0.2$. The hidden dimension size is set to $64$ for all hidden layers. The model is trained for $50$ epochs by minimizing the cross-entropy loss function with the Adam optimizer and a learning rate of $0.001$. At the end of each epoch, we compute the performance of the model on the validation set, and we choose as our final model the one that achieved the smallest loss on the validation set.
@@
Figure 7 illustrates the Wasserstein distance with $p=1$ between groups when the EMD (Top Right) and the Hausdorff distance (Bottom Right) are used as ground metrics. It also shows the drop in accuracy when the $\textsc{NN}_{\textsc{mean}}$ (Top Left) and $\textsc{NN}_{\textsc{max}}$ (Bottom Left) models are trained on one domain and evaluated on the others. Each row corresponds to one specific model, e. g., the first row represents the model trained on the reviews for books. We observe that in general the drop in accuracy follows a similar pattern with the distance between the source and target domains, i. e., the higher the distance, the higher the drop in accuracy. We also computed the Pearson correlation between the drop in accuracy and the domain dicrepancies. We found that the Wasserstein distance based on EMD highly correlates with the accuracy drop of $\textsc{NN}_{\textsc{mean}}$ ($r=0.917$), while there is an even higher correlation between the Wasserstein distance based on Hausdorff distance and the accuracy drop of $\textsc{NN}_{\textsc{max}}$ ($r=0.941$).
E.5 Predictive Performance of Neural Networks for Sets
We next evaluate the $\textsc{NN}_{\textsc{sum}}$, $\textsc{NN}_{\textsc{mean}}$ and $\textsc{NN}_{\textsc{max}}$ models on four classification and regression datasets, namely ModelNet40, Polarity, IMDB and IMDB-BINARY. The first two datasets are described in section 4. IMDB contains movie reviews from the IMDb database (Maas et al., 2011). The targets are the ratings that accompany the reviews ($10$ different values). We treat this task as a regression problem. IMDB-BINARY is a standard graph classification dataset (Yanardag and Vishwanathan, 2015), commonly used for evaluating graph kernels and graph neural networks. Each graph of the IMDB-BINARY dataset was represented as a multiset of the degrees of its nodes. The results are illustrated in Table 3. Each experiment was repeated $5$ times with different random seeds, and for each dataset we report average accuracy (for ModelNet40, Polarity and IMDB-BINARY) or average root mean square error (for IMDB) on the dataset’s test set and the corresponding standard deviation.
We can see that $\textsc{NN}_{\textsc{max}}$ outperforms the other models on ModelNet40. A possible explanation is that all input multisets have the same size, and in such a setting the max aggregator is Lipschitz continuous with respect to all three considered distance functions. Therefore, $\textsc{NN}_{\textsc{max}}$ can effectively capture the distances between the point clouds in ModelNet40.
Polarity consists of short reviews (average number of terms = $20$). Therefore, whether a review is positive or negative depends primarily on the presence of one or a few terms that indicate sentiment. These terms can be considered extreme elements, and the Hausdorff distance relies on such extreme elements when comparing its inputs. This distance function thus seems to be suitable for this task and potentially explains why $\textsc{NN}_{\textsc{max}}$ is the best-performing model on this dataset.
Table 3: Average performance (accuracy or root mean square error) of the $\textsc{NN}_{\textsc{sum}}$, $\textsc{NN}_{\textsc{mean}}$ and $\textsc{NN}_{\textsc{max}}$ on the four benchmark datasets.
ModelNet40 $(\uparrow)$ Polarity $(\uparrow)$ IMDB $(\downarrow)$ IMDB-BINARY $(\uparrow)$
$\textsc{NN}_{\textsc{sum}}$ 60.07 $\pm$ 2.12 76.93 $\pm$ 2.42 0.2210 $\pm$ 0.0085 70.20 $\pm$ 1.72
$\textsc{NN}_{\textsc{mean}}$ 63.41 $\pm$ 0.98 77.11 $\pm$ 2.34 0.2159 $\pm$ 0.0033 67.00 $\pm$ 2.52
$\textsc{NN}_{\textsc{max}}$ 77.21 $\pm$ 1.09 78.14 $\pm$ 1.94 0.2259 $\pm$ 0.0017 61.40 $\pm$ 2.57
The reviews contained in the IMDB dataset are much longer than those in the Polarity dataset (average number of terms = $254$). Due to the potential presence of outlier terms, the Hausdorff distance may not accurately capture document similarity. The matching distance can also be sensitive to document length. In contrast, EMD captures the overall semantic alignment between documents, and compares documents based on their overall meaning. This makes EMD more suitable for this task, which is empirically confirmed by the superior performance of $\textsc{NN}_{\textsc{mean}}$ compared to the other two models.
$\textsc{NN}_{\textsc{sum}}$ is the best-performing method on the IMDB-BINARY dataset. In this dataset, capturing both the number of nodes and their degrees is essential. The matching distance is well suited for this task, and the stronger performance of the $\textsc{NN}_{\textsc{sum}}$ model, which is Lipschitz continuous with respect to this distance (under certain conditions), supports this intuition.